Suppose x has a distribution with μ = 13 and σ = 6. (a) If a random sample of size n = 35 is drawn, find μx, σ x and P(13 ≤ x ≤ 15). (Round σx to two decimal places and the probability to four decimal places.) μx =

σx =

P(13 ≤ x ≤ 15) =

(b) If a random sample of size n = 61 is drawn, find μx, σ x and P(13 ≤ x ≤ 15). (Round σ x to two decimal places and the probability to four decimal places.)

μx =

σ x =

P(13 ≤ x ≤ 15) =
Answer #1

(Date Solution in given that, M = 13 a n=35 Mx = M = 13 6x = 6 P(13<X L15)=P/13-13 X-4X 15-13) (1.01 1.01 =pCOLZL 1.98) = P(Z